Circular No. 6836

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998U IN ANONYMOUS GALAXY N. Regnault and O. Perdereau, on behalf of the EROS collaboration

(cf. IAUC 6760, 6762), report the discovery of an apparent supernova (V ∼ 20) on Mar. 3 (and confirmed on Mar. 4) on images taken with the Marly 1-m telescope at the European Southern Observatory (ESO). SN 1998U is located at α = 11^{h}36^{m}32^{s}.20, δ = −12^{o}07′26′′.7 (equinox 2000.0), which is 2′′.4 west and 4′′.2 north of the host galaxy. Additional images of SN 1998U were obtained on Mar. 6 at the ESO 1.5-m Danish telescope. SN 1998U was not present on a CCD frame taken on 1997 Feb. 2 (limiting mag 22).

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD A. Crotts, Columbia University; and S. Heathcote, Cerro Tololo In-

teramerican Observatory (CTIO), report on data obtained with the CTIO 4-m telescope on 1996 Dec. 5 and 1998 Feb. 17 UT (échelle spectrograph), and on 1997 Sept. 23 (RC spectrograph): “The increase in strength of the blue wing of the Hα emission line, as reported on IAUC 6831, is difficult to confirm with certainty in our data. Some weak emission lines have formed in the near-infrared portion of the spectrum, particularly [Fe II] at 768.6, 889.1, and 905.1 nm, and appear offset to the northeast of the supernova by ∼ 0′′.5 (for the slit at p.a. 50^{o}). (Stronger lines due to [Fe II] at 715.5, 717.2, 738.8, 745.2, and 861.7 nm are detected also in RC spectrograph data obtained on 1997 Jan. 15 by J. Maza and M. Phillips, while [Fe II] 768.6, 889.1, and 905.1 nm would have been detected at the strength found on 1997 Sept. 23, but were not.) Previously detected lines over the range 570–1080 nm are the H I Paschen series (all terms 7 to 17); [Fe VII] 608.7 nm; [Ar III] 713.5 and 775.1 nm; [Ar V] 700.5 nm; [Cl II] 857.9 nm; and [Ca II] 729.1 and 732.3 nm; [Ni II] 737.8 nm; [Cr II] 812.5, 822.9, and 830.8 nm; and perhaps [C I] 984.9 nm and [Ar IV] 723.7 and 726.2 nm. Also seen are the usual strong lines of [N II], He I, [O I], [O II], [S II], [S III] and Hα. New, unidentified, very weak lines are also seen at rest wavelengths 704.56, 710.37, and 711.98 nm (with a typical accuracy of ± 0.02 nm and calculated for the supernova’s redshift of 289 km/s). All of the above lines have narrow components with velocity widths less than ∼ 100 km/s. This spectrum appears to be dominated by low-density, low-ionization shock excitation, especially to the northeast of the supernova (consistent in position with the previously reported ‘hot spot’; cf. IAUC 6665, 6710).”

1998 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green